# Multimodal classification of forest biodiversity potential from 2D orthophotos and 3D airborne laser scanning point clouds

### v2

@@ 4.2.1 Near-infrared for forest biodiversity assessment
ResNet 1 ResNet 2 77.9% 77.5% 79.5% ± 0.6 +1.8 +1.6
ResNet 2 ResNet 3 77.5% 81.2% 79.9% ± 0.6 +0.5 -1.3
ResNet 3 ResNet 1 81.2% 77.9% 80.9% ± 0.2 +1.4 -0.3
2D-2D Average 78.9% 78.9% 80.1% +1.2 +0.0
PointVector 1 PointVector 2 75.6% 76.1% 74.9% ± 2.4 -0.9 -1.2
PointVector 2 PointVector 3 76.1% 76.2% 76.2% ± 1.0 +0.1 +0.0
PointVector 3 PointVector 1 76.2% 75.6% 75.8% ± 1.2 -0.1 -0.4
3D-3D Average 76.0% 76.0% 75.6% -0.3 -0.5
ResNet 1 PointVector 1 77.9% 75.6% 81.4% ± 0.5 +4.7 +3.5
ResNet 2 PointVector 2 77.5% 76.1% 81.2% ± 0.2 +4.4 +3.7
ResNet 3 PointVector 3 81.2% 76.2% 82.7% ± 1.2 +4.0 +1.5
2D-3D Average 78.9% 76.0% 81.8% +4.4 +2.9
Table 7: Multimodal fusion performance comparison. Analysis of model fusion performance between different modality pairs (2D-2D, 3D-3D, and 2D-3D), showing overall accuracy for individual models and fused versions along with performance gains relative to average and maximum single model accuracies.
We observed that fusion of feature embeddings from 2D orthophotos and 3D ALS point clouds leads to consistent performance gains, with a mean gain (avg) of 4.4 percentage point improvement. 3D + 3D fusion does not show synergy, indicating that different instances of the PointVector model extracts consistent feature embeddings. 2D + 2D fusion performs marginally better than individual models, indicating that different instances of trained ResNet models actually capture complementary features, while not as much as observed for 2D + 3D MLP fusion.
Figure 11: Low forest biodiversity potential example. Predictions and as associated confidence scores from from all tested methods for an area with low forest biodiversity potential (low_2019_24). The 2D + 3D MLP Fusion methods achieve the best performance with a mean accuracy of 84.0%. Accuracies for each model are shown in the top right of each prediction image.
@@ 4.6 Qualitative analysis
To better understand the tested methods’ performance and behavior, we analyzed their predictions on two representative forest areas with contrasting biodiversity potential in Figures 11 and 12. Figure 11 shows predictions on a low forest biodiversity potential area (id: low_2019_24) from the BioVista test set, containing 25 samples. Both MLP Fusion methods (end-to-end trained and trained on pre-computed feature embeddings) notably outperforms single-modality methods, achieving an accuracy of 84.0% compared to ResNet’s 36.0% and PointVector’s 0.0%. The example highlights instances where MLP Fusion methods manages to correctly classify the target class, even when both single-modality methods fail to do so. Additionally, we observe that the end-to-end and pre-computed variants of MLP fusion generally yield similar predictions with minor confidence score variations.
Figure 12: High forest biodiversity potential example. Predictions and as associated confidence scores from all tested methods on a high forest biodiversity potential area (high_2021_24) from the BioVista test set, except 2D + 3D MLP Fusion (end-to-end). Accuracies for each model are shown in the top right of each prediction image.
@@ 5 Discussion
Our work demonstrates that remote sensing combined with deep learning offers a promising pathway for large-scale forest biodiversity monitoring, though important limitations and considerations must be acknowledged.
@@ 5.1 Implications
Our results, shown in Table 6 and Figure 10 demonstrate that deep learning models can effectively assess forest biodiversity potential on a coarse-grained level using remote sensing data, with our best model achieving an overall accuracy of 81.4%. Particularly noteworthy is how combining 2D orthophotos and 3D ALS point clouds, through deep learning-based multimodal fusion, enhances classification performance. This performance enhancement shows that spectral and structural information complement each other in meaningful ways for forest biodiversity assessment, as clearly demonstrated in Section 4.5.
@@ 5.2 Limitations
Another limitation pertains to the seasonality of the orthophoto and ALS point cloud data. As noted in Section 2.3, the images were acquired between March 1st and May 1st, corresponding to the leaf-off period for many deciduous species in Danish temperate forests. We acknowledge that forest canopy structure and spectral responses vary significantly across seasons. Utilizing orthophotos and ALS point clouds captured during the summer or autumn would provide different spectral information related to chlorophyll content, vegetation health, and potentially allow for better or worse differentiation based on leaf characteristics. This seasonal variation represents a form of temporal or concept drift, which is a known challenge for machine learning models operating on environmental data over time (Nikolov et al., 2021).
@@ 5.3 Future work
A crucial next step is to evaluate the generalizability of our approach across different forest biomes. While our method shows promise for temperate forests in Denmark, tropical, boreal, and other forest types present distinct challenges and characteristics. Tropical forests, for instance, typically exhibit much higher species diversity and more complex canopy structures than temperate forests, which could affect both the ALS point cloud patterns and spectral signatures in orthophotos. While our current binary classification approach provides a solid foundation for forest biodiversity assessment, future work should explore more fine-grained classification schemes. Rather than categorizing forests as having high- or low biodiversity potential, a multi-class approach could distinguish between several levels of biodiversity potential. This would provide more nuanced insights for conservation prioritization and forest management.
@@ 6 Conclusion
This study demonstrates the potential of 2D orthophotos, 3D ALS point clouds and a combination of the two, for assessment of forest biodiversity potential using deep learning methods. Through evaluation using our introduced BioVista dataset, we have shown that deep learning methods can successfully classify forest biodiversity potential, with our best fusion model achieving an overall accuracy of 81.4%. The superior performance achieved through multimodal fusion suggests that combining structural and spectral information provides a more comprehensive and reliable assessment of forest biodiversity. This multimodal potential should be tested for predicting other forest-related variables as well. This research represents a step towards developing methods that can support forest biodiversity monitoring and conservation efforts, but further improvements in accuracy and robustness are needed before these methods could be considered for practical applications. Future work should incorporate a more fine-grained classification scheme and test the methodology across different forest biomes.
@@ Appendix A Forest area performance
We analyzed classification results on a selections of forest areas in the BioVista test set. The results are seen in Table 8. The results reveal considerable variation in model performance across different areas. For some areas, all tested models display perfect classification accuracy, while others prove challenging for some or all methods.
@@
We found similar patterns of variation in low forest biodiversity areas, where several areas (e.g., low_2020_31, low_2023_41) were classified perfectly by all models, while others (e.g., low_2023_29) proved particularly challenging with near-zero accuracy across all approaches. The MLP Fusion model showed notably strong performance on moderately difficult low forest biodiversity areas, often achieving the highest accuracy among all models.

### v3

@@ 4.2.1 Near-infrared for forest biodiversity assessment
2D-2D
ResNet 1 ResNet 2 78.3 % 77.8 % 79.0 % +1.0 +0.8
ResNet 2 ResNet 3 77.8 % 81.3 % 80.5 % +0.9 -0.8
ResNet 3 ResNet 1 81.3 % 78.3 % 80.9 % +1.1 -0.5
Average 79.1 % 79.1 % 80.1 % +1.0 +1.0
3D-3D
PointVector 1 PointVector 2 76.0 % 76.0 % 76.1 % +0.1 +0.1
PointVector 2 PointVector 3 76.0 % 75.9 % 76.0 % +0.0 -0.0
PointVector 3 PointVector 1 75.9 % 76.0 % 76.4 % +0.5 +0.4
Average 76.0 % 76.0 % 76.2 % +0.2 +0.2
2D-3D
ResNet 1 PointVector 1 78.3 % 76.0 % 81.3 % +4.1 +3.0
ResNet 2 PointVector 2 77.8 % 76.0 % 80.2 % +3.3 +2.4
ResNet 3 PointVector 3 81.3 % 75.9 % 81.7 % +3.1 +0.3
@Average 79.1 % 76.0 % 81.1 % +3.5 +1.9
We observed that ensembling models from 2D orthophotos and 3D ALS point clouds leads to consistent performance gains, with a mean gain (avg) of 3.5 percentage point improvement. 3D + 3D does not show strong synergy, indicating that different instances of the PointVector model extracts consistent feature embeddings. 2D + 2D performs marginally better than individual models, indicating that different instances of trained ResNet models actually capture complementary features, while not as much as observed for 2D + 3D.
Figure 10: Qualitative analysis – low forest biodiversity potential example. Predictions and as associated confidence scores from from all tested methods for an area with low forest biodiversity potential (low_2019_24). The 2D + 3D MLP Fusion methods achieve the best performance with a mean accuracy of 84.0%. Accuracies for each model are shown in the top right of each prediction image.
@@ 4.6 Qualitative analysis
To better understand the tested methods’ performance and behavior, we analyzed their predictions on two representative forest areas with contrasting biodiversity potential in Figures 10 and 11. Figure 10 shows predictions on a low forest biodiversity potential area (id: low_2019_24) from the BioVista test set, containing 25 samples. Both MLP Fusion methods (end-to-end trained and trained with frozen backbones) notably outperformed single-modality methods, achieving an accuracy of 84.0% compared to ResNet’s 36.0% and PointVector’s 0.0%. The example highlights instances where MLP Fusion methods manage to correctly classify the target class, even when both single-modality methods fail to do so. Additionally, we observed that the end-to-end and frozen backbone variants of MLP fusion generally yielded similar predictions with minor confidence score variations.
Figure 11: Qualitative analysis – high forest biodiversity potential example. Predictions and as associated confidence scores from all tested methods on a high forest biodiversity potential area (high_2021_24) from the BioVista test set. Accuracies for each model are shown in the top right of each prediction image.
@@ 5 Discussion
Our work demonstrates that remote sensing combined with deep learning offers a promising pathway for large-scale forest biodiversity potential monitoring, though important limitations and considerations must be acknowledged.
@@ 5.1 Implications
A central conclusion of our work is that 2D- and 3D multimodal fusion provides a significant performance advantage against single modality methods. Our results clearly show that models fusing 2D orthophotos and 3D ALS point clouds consistently outperform models based on a single modality (see Table 6). Our analysis in Section 4.5 further demonstrates that this is not merely an ensembling effect – the performance gain stems from the fact that the spectral information from orthophotos and the structural information from ALS point clouds provide distinct and complementary features for assessing biodiversity potential. Ensembling two models of the same type (e.g., 2D + 2D) yields only marginal gains compared to the substantial improvements seen with 2D + 3D fusion (see Table 8).
Our results, shown in Table 6 and Figure 9, demonstrate that deep learning models can effectively assess forest biodiversity potential on a coarse-grained level using remote sensing data, with our best model achieving an overall accuracy of 82.0% when separating low and high biodiversity potential areas.
Our results in Table 7 justify the binarization of biodiversity potential, as experiments with three classes showed reduced performance and higher confusion between intermediate levels. This outcome suggests that the current combination of remote sensing data and the available biodiversity potential definitions may not yet be sufficient to capture finer gradients reliably. Despite this simplification, binary predictions of biodiversity potential remain valuable for identifying and prioritizing areas for field surveys, conservation planning, and targeted data collection. While more detailed and harmonized biodiversity potential metrics would be desirable, such scalable indicators are not yet available across Europe or Scandinavia. The current HNV framework, though imperfect, serves as a practical and reproducible proxy for large-scale biodiversity assessments and provides a solid foundation for future refinement as improved data and indicators emerge.
@@ 5.2 Limitations
Another limitation pertains to the seasonality of the orthophoto and ALS point cloud data. As noted in Section 2.3, the images were acquired between March 1st and May 1st, corresponding to the leaf-off period for many deciduous species in Danish temperate forests. We acknowledge that forest canopy structure and spectral responses vary significantly across seasons. Utilizing orthophotos and ALS point clouds captured during the summer or autumn would provide different spectral information related to chlorophyll content, vegetation health, and potentially allow for better or worse differentiation based on leaf characteristics. This seasonal variation represents a form of temporal or concept drift, which is a known challenge for machine learning models operating on environmental data over time (nikolov-2021-seasons).
@@ 5.3 Future work
A crucial next step is to evaluate the generalizability of our approach across different forest biomes. While our method shows promise for temperate forests in Denmark, tropical, boreal, and other forest types present distinct challenges and characteristics. Tropical forests, for instance, typically exhibit much higher species diversity and more complex canopy structures than temperate forests, which could affect both the ALS point cloud patterns and spectral signatures in orthophotos. Although the model is robust to terrain variation, incorporating absolute elevation as an additional input could be beneficial in more topographically complex regions, since it currently only learns from plot-relative height information. While our binary classification approach provides a proof of concept for forest biodiversity potential assessment, future work should explore more fine-grained classification schemes. A multi-class approach that reliably distinguishes between several levels of biodiversity would provide more nuanced insights for conservation prioritization and forest management.
@@ 6 Conclusion
This study demonstrates the potential of 2D orthophotos, 3D ALS point clouds and a combination of the two, for assessment of forest biodiversity potential using deep learning methods. Through evaluation using our introduced BioVista dataset, we have shown that deep learning methods can successfully classify forest biodiversity potential, with our best fusion model achieving an overall accuracy of 82.0% when separating low and high potential areas. The superior performance achieved through multimodal fusion suggests that combining structural and spectral information provides a more comprehensive and reliable assessment of forest biodiversity. This multimodal potential should be tested for predicting other forest-related variables as well. This research represents a step towards developing methods that can support forest biodiversity monitoring and conservation efforts, but further improvements in accuracy, granularity, and robustness are needed before these methods could be considered for practical applications. Future work should incorporate a more fine-grained classification scheme and test the methodology across different forest biomes.
@@ Appendix A Forest area performance
We analyzed binary classification results on a selection of forest areas in the BioVista test set. The results are shown in Table 9. The results reveal considerable variation in model performance across different areas. For some areas, all tested models display perfect classification accuracy, while others prove challenging for some or all methods.
@@
We found similar patterns of variation in low forest biodiversity potential areas, where several areas (e.g., low_2020_31, low_2023_41) were classified perfectly by all models, while others (e.g., low_2023_29) proved particularly challenging with near-zero accuracy across all approaches. The MLP Fusion model showed notably strong performance on moderately difficult low forest biodiversity potential areas, often achieving the highest accuracy among all models.
@@
Appendix B Intermediate Fusion with Short Cuts
In addition to the 2D and 3D fusion approaches presented in Section 3 (late fusion through confidence-based ensembling and feature-level concatenation), we carried our an additional fusion experiments using intermediate fusion with short cuts. The approach is illustrated in Figure 13.
Figure 13: Intermediate fusion of 2D orthophotos and 3D ALS point clouds with short cuts. Feature embeddings from intermediate layers in the ResNet and PointVector backbones are pooled and concatenated into the fully connected layers of the MLP head, using short cuts.
We train the network for 40 epochs, using the same hyperparameters as described in Table 5 and a learning rate of $10^{-5}$ for the ResNet and PointVector backbones, similar to our MLP Fusion (end-to-end). In order to achieve a batch size of 64 we use gradient accumulation (8 mini batches of batch size 8).
We implemented the intermediate fusion with short cuts, by pooling (average pool) feature embeddings from intermediate layers in the ResNet and PointVector backbones and concatenating them into the fully connected layers of the MLP fusion head. The size of the hidden layers of the MLP being 1024, 1024, 512, and 256. We used pre-trained weights for ResNet and PointVector, obtained through the single modality experiments described in Section 4.2.1 and 4.3.
The intermediate fusion approach is evaluated on the BioVista dataset and compared to the other tested 2D and 3D fusion approaches in Table 10. The 2D + 3D Intermediate Fusion with shortcuts approach, while hypothesized to benefit from features at multiple abstraction levels, did not yield performance improvements over other tested fusion techniques. In terms of overall accuracy (79.7% OAcc) and mean accuracy (79.8% MAcc), it ranked lower than ensembling and both MLP fusion strategies. While its performance on the high forest biodiversity potential class was competitive with MLP fusion methods, it was surpassed by ensembling. For the low forest biodiversity potential class, it performed better than ensembling but was weaker than the MLP fusion approaches. This approach did, however, still offer an improvement over using single modalities, which achieved OAccs of 76.7% ± 2.6 and 75.8% ± 0.4 for strictly 2D and 3D respectively. The observed suboptimal performance could be attributed to several factors. The use of average pooling to aggregate intermediate features might dilute crucial discriminative information. Exploring alternatives like max pooling could be beneficial. Furthermore, the increased complexity from concatenating features from multiple layers might pose challenges for the MLP head to effectively learn optimal representations, potentially requiring more sophisticated architectural choices or regularization for the fusion module itself.
Table 10: Performance comparison of fusion approaches. We extend Table 6 with the results from the intermediate fusion with short cuts to get an overview of the performance of all the tested fusion approaches. All metrics are means with associated standard deviation across seven runs with the same hyperparameter settings.
Model OAcc MAcc $\textbf{Acc}_{\text{high}}$ $\textbf{Acc}_{\text{low}}$
2D + 3D Ensembling 80.5% ± 2.0 80.6% ± 1.9 89.5% ± 2.7 71.8% ± 6.0
2D + 3D MLP Fusion (frozen backbones) 81.4% ± 0.5 81.3% ± 0.5 85.8% ± 4.5 76.9% ± 3.8
2D + 3D Intermediate Fusion 79.7% ± 1.2 79.8% ± 1.1 86.2% ± 3.5 73.4 % ± 5.2
3 List of figure captions